Circular No. 6997

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998dk IN UGC 139 A. V. Filippenko, University of California at Berkeley, reports that in-

spection of an uncalibrated CCD spectrum of SN 1998dk (cf. IAUC 6991) obtained on Aug. 24 by C. De Breuck (Lawrence Livermore National Laboratory) with the 3-m Shane reflector at Lick Observatory reveals that the object is a type-Ia supernova close to maximum brightness, with strong Si II lines.

SUPERNOVA 1998dm IN MCG –01-4-44 Filippenko also notes that inspection of an uncalibrated CCD spectrum

of SN 1998dm (cf. IAUC 6993), obtained as above by De Breuck, reveals that this object is a type-Ia supernova prior to maximum brightness. The spectrum is unusually red, and the Si II line near 575 nm seems strong relative to Si II 615-nm, suggesting that this is a subluminous type-Ia event.

SAX J1748.9–2021, MX 1746–20, AND NGC 6440 J. in ’t Zand and J. Heise, Space Research Organization of the Nether-

lands (SRON), Utrecht; A. Bazzano and P. Ubertini, Istituto di Astrofisica Spaziale, CNR, Rome; M. J. S. Smith, BeppoSAX Science Operation Center (BSOC), Rome, and SRON; J. M. Muller, BeppoSAX Science Data Center, Rome, and SRON; and V. Torroni, BSOC, communicate: “During an observation with the Wide Field Cameras on BeppoSAX on Aug. 22.48– 23.19 UT, an x-ray transient and burster was detected at α = 17^{h}48^{m}52^{s}, δ = −20^{o}21′.4 (equinox 2000.0; 99-percent confidence error radius 2′). We designate the source SAX J1748.9–2021. The centroid is 0′.3 from the center of the globular cluster NGC 6440, which makes an association very likely. The average intensities are ∼ 30 mCrab in 2–10 keV and ∼ 45 mCrab in 10–25 keV. On an hourly basis, the source does not show variability above an upper limit of 20 percent. One x-ray burst was detected from the same transient, with a peak flux of ∼ 0.7 Crab (2–25 keV), indicating that the transient is a low-mass x-ray binary. This is the first time that an x-ray transient was seen coincident with NGC 6440 since December 1971, when MX 1746–20 was discovered (Markert et al. 1975, Nature 257, 32), which was active for 1.5–3 months with a peak flux of ∼ 150 mCrab and had a large 90-percent confidence error box of ∼ 6′ by 30′ (Forman et al. 1976, Ap.J. 207, L25). Follow-up observations with the narrow-field instruments on BeppoSAX and with RXTE have been scheduled.”

1998 August 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green